Grade 11Inorganic Chemistry

All formulas of atomic structure. Some formulas is missing.convesion of R in m and cm

Ayushi
7 Years agoGrade 11

1 Answer

Shivangi Khatter
7 Years ago
radius = 0.529 * n2/ZÅ
=0.529 * 10-10* n2/Z m
=0.529 * 10-8 * n2/Zcm